# Investigating the correlation between ZTF TDEs and IceCube high-energy neutrinos

### v1

@@ 3 Analysis Method
We perform a systematic search in our samples based on the spatial and temporal criteria. The initial step involves examining the spatial correlation between TDEs and alert events. This is realized by cross-correlating the coordinates of TDEs within of the alert events. In this step, we search for spatially associated events, using the alert events sample with 30 deg2 and the whole alert events sample, respectively. According to the of IceCube alert events, the calculation of for alert events is given by Kouch et al. (2024),
where, and are the four directional errors, respectively. In this case, we employ the of alert events in these four directions.
In the next step, we identify the temporal coincident events from the spatially correlated events, i.e., these TDEs that were discovered before the arrival dates of alert events. Therefore, the test-statistic (TS) parameter is the time delay (, the arrival date of the alert event subtracting the discovery date of TDE). If one TDE is found within the of an alert event and 0, the TS value would be increased by 1. During the calculation of TS values, we allow that a single TDE or alert event can contribute repeatedly to the TS values. Specifically, if one alert event is spatially and temporally associated with TDEs, or if one TDE is spatially and temporally associated with alert events, they can each contribute a TS value of (i.e., the total TS will be plus ).
As mentioned above, we consider two methods to handle the sample of alert events. The first method involves excluding alert events of 30 deg2 (hereafter referred to CUT method). In this scenario, the number of TDEs that are spatially-temporally associated with the alert events corresponds to the TS values.
The other method includes the whole alert events sample while assigning a weight factor that combines the and signalness of alert events (hereafter referred to WEIGHT method). The weighting factor can be expressed as (Kouch et al., 2024),
@@
where, and are the signalness and for the -th alert event, respectively. is the median of of alert events sample. In the weighting scheme, the weight factor is implemented in the calculation of TS values by multiplying when TDEs are found to be spatially-temporally associated with the -th alert event.
@@
According to the of the alert events and Eq. 2, we estimate that the of our alert events sample is 6.88 deg2. We present the TS values using the heating map after considering the weight factor, as shown in Fig 3, where the color bar is the TS value when one TDE is found to be associated with an alert event, the black dashed line indicates the median value of our alert events sample (6.88 deg2), inverted triangles are alert events that are not associated any TDE and the red stars present alert events that are spatially-temporally associated with at least one TDE, respectively. One can see that the values of alert events that are associated with TDEs are larger than the median value.
Based on these two methods for handling the alert events sample, three MC simulation methods are adopted to perform a correlation analysis between the TDEs population and the alert events population. They are described as follows.
@@
Method I: similar to previous works (e.g., Plavin et al., 2020; Hovatta et al., 2021), due to the sensitivity of IceCube only depends on the zenith angle (Aartsen et al., 2017d), we only randomize the RA of the alert events, while keeping the DEC fixed. The and arrival date are consistent with their observed one for each alert event, and the information about the TDEs remains fixed.
@@
Method II: we not only randomize the RA of alert events, as in Method I, but also simulate the TDE sample using the method described in Lu et al. (2024). The method for simulating the TDE sample is as follows. We randomly select and from the observed lists of RA and DEC, respectively, to represent a mock source. Furthermore, we impose an additional random shift to within a region of radius. For each source, we ensure that it remains within the same sky region as the ZTF-BTS sample, which covers the northern sky with a DEC range from to 90∘, while excluding the Galactic plane region (). When we obtain the number of mock sources as the same as the observed one, we use the KS test to ensure that the mock sources preserve the distribution pattern of the actual sources, i.e., and of mock sources are performed KS test, respectively.
@@
Method III: the same as Method II, we also randomize the RA of alert events and TDE positions. The method used to simulate the TDE sample is consistent with Buson et al. (2022, 2023); Bellenghi et al. (2023). We simulate TDEs by randomizing the positions of the sources within a 10∘-radius range from their original positions. During this process, we also ensure that each pseudo-source remains within the sky region of the ZTF-BTS sample. The KS test is also performed in the same manner as in Method II.
@@
By realizing 104 times MC simulations to obtain the TS distribution. We can then estimate the chance probability (i.e., the value of rejecting the null hypothesis) by counting the fraction of TS values that are equal to or larger than the observed value,
@@
where is the number of TS values equal to or larger than the observed one and is the total number of MC simulations (Davison & Hinkley, 1997). Then, we can estimate the values for these three MC simulation methods.
@@ 4 Results
We first systematically search for associated events that satisfy spatial and temporal criteria. After adopting the CUT method, there are 122 alert events remaining in our sample. Three associated events are found, and two of these events have IR flares, which are listed in Table 2. Apart from AT2019dsg, two other TDEs are associated with two bronze alert events.
Under the WEIGHT method, all 143 alert events are utilized to search for associated events. We found that there are seven occurrences of TDEs within the of five alert events. Among them, one TDE (AT2019azh) is located within of two alert events. Additionally, the of two alert events each includes two TDEs. However, we note that the values of these two alert events are significantly large ( 10∘), contributing only negligible TS values. In total, six distinct TDEs are found within of five alert events, two of which have IR flares, as listed in Table 3.
RAalert | DECalert | Signalness | TDE event | TS1 valueb | IR flarec | |||
(deg) | (deg) | (deg) | (days) | |||||
124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 1 | Yes |
47.20 | -3.28 | 2.1 | 22/05/24 | 0.336 | AT2020afhd | 581 | 1 | No |
314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1 | Yes |
@@
c This column indicates whether the TDEs have IR flares.
RAalert | DECalert | Datealert | Signalness | TDE event | TS1 valueb | IR flarec | ||
(deg) | (deg) | (deg) | (days) | |||||
239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020vwl | 1244 | 0.0056 | No |
239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020pj | 1526 | 0.0056 | No |
127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2022bdw | 522 | 0.010 | No |
127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2019azh | 1596 | 0.010 | Yes |
124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 0.17 | Yes |
47.20 | -3.28 | 2.1 | 22/05/24 | 0.336 | AT2020afhd | 581 | 0.17 | No |
314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 0.15 | Yes |
@@
b The 1-event TS represents the value that an association event can contribute to the total TS of the correlation.
c This column indicates whether the TDEs have IR flares.
@@
CUT method | 0.16 | 0.15 | 0.15 | All TDEs |
WEIGHT method | 0.34 | 0.32 | 0.32 | All TDEs |
CUT method | 0.027 | 0.015 | 0.020 | TDEs with IR flare |
WEIGHT method | 0.044 | 0.044 | 0.047 | TDEs with IR flare |
By realizing 104 times MC simulations, the results obtained from the three MC methods are summarized in Table 4, which includes the results from both the CUT method and the WEIGHT method. Using the all TDEs sample, the best -values are 0.15 and 0.32 for the CUT method and WEIGHT method, respectively. When using the subsample of TDEs with IR flares, the best -values are 0.015 and 0.044 for the CUT method and WEIGHT method, corresponding to 2.43 and 2.01, respectively. The most significant correlation with neutrinos is obtained from the subsample of TDEs with IR flare at a confidence level, which may indicate that TDEs with IR flare are more likely to correlate with IceCube neutrinos compared to all TDEs sample.
@@ 5.1 The systematic errors of alert events
Although the IceCube collaboration strives to consider all potential errors in estimating the size of the high-energy neutrino error regions, there may still be some unknown systematic errors, such as those arising from ice inhomogeneities, that cannot be completely addressed (e.g. Abbasi et al., 2023b). An upper limit of 1∘ is estimated for this systematic error by Aartsen et al. (2013c). While efforts to constrain this parameter are ongoing (Abbasi et al., 2021b, 2022c), the most recent findings suggest that the associated error is likely negligible (Abbasi et al., 2022c, 2023b).
In this work, we add a systematic error upper limit of 1∘ in quadrature to the four directions of RA and DEC, so that the final uncertainty in each direction is . We then search for associated events among all alert events (having added 1∘ systematic error) and TDEs. However, no additional associated events are found when compared to those listed in Table 3. Therefore, the value is expected to increase as the systematic error grows larger within the range of . When a 1∘ systematic error is added to each alert event, the of the alert events sample is 10.02 deg2. We only adopt the WEIGHT method to perform the effect of systemic error on the estimation of values. The values estimated by three MC simulation methods are listed in Table 5, showing an increase of to compared to those in Table 4.
@@
WEIGHT+Syst. | 0.44 | 0.45 | 0.43 | All TDEs |
WEIGHT+Syst. | 0.060 | 0.054 | 0.068 | TDEs with IR flare |
WEIGHT+Delay | 0.29 | 0.30 | 0.28 | All TDEs |
WEIGHT+Delay | 0.011 | 0.043 | 0.032 | TDEs with IR flare |
@@ 5.2 Time delay factor
In this analysis, the temporal coincidence criterion is required that the discovery dates of TDEs are earlier than the arrival dates of alert events. In this scenario, we estimate the values of correlation between TDEs/subsample of ZTF-BTS catalog and IceCube neutrinos, and the most significant correlation is given by the TDEs with IR flares subsample.
@@
Therefore, we tentatively impose a time delay factor into the weighting scheme to reduce the weight of associated events with large time delays. By combining an additional time delay factor into the weight factor of Eq. 3a, 3b, we reduce the impact of events with large time delays, i.e.
@@
After adding the time delay factor, the TS values are mainly contributed by the AT2019dsg event. By realizing 104 times MC simulations, the values for the three MC simulation methods are summarized in Table 5. Compared to the results without the time delay factor, the best value of the all-TDEs sample decreases to 0.28, while that for the TDEs with IR flares subsample decreases to 0.011 (). After introducing the time delay factor, the subsample with IR flare also presents the most significant correlation with IceCube neutrinos compared to the all-TDEs sample.
@@ 6 Summary
Currently, the main contributors to the IceCube neutrino flux are still obscure. Among various astrophysical sources, TDEs have been suggested as possible emitters of IceCube neutrinos. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether TDEs could really serve as neutrino emitters. In this work, we study the correlation between the TDEs/TDE candidates from the ZTF-BTS catalog and the population of IceCube neutrinos. Considering that good spatial localization is crucial for establishing reliable association, we employ two methods to handle the alert events with large error circles. One method involves directly excluding alert events with error area 30 deg2, while another method introduces a weighting factor. We quantify the significance of the correlation using three MC simulation methods. By performing MC simulations, we find that the subsample of TDEs with IR flares presents the most significant correlation with IceCube neutrinos, though still below the confidence level. We also discuss the case that includes a systematic error in the events’ error radii. Given that only the statistical error is publicly available for alert events, we add a systematic error of , which is the upper limit of the systematic error estimated by Aartsen et al. (2013c), in quadrature to the statistical error. There are no additional associated events arising from the inclusion of systematic error and the obtained results could be treated as the upper limits of the values. Furthermore, for the potentially associated events reported so far, the arrival dates of neutrinos are within an interval of year after the discovery dates of the TDEs/TDE candidates. Therefore, we tentatively introduce a time delay () term into the weighting factor in our MC simulation to reduce the weight of large time-delayed events. By combining the time delay factor, the best values decrease to 0.28 and 0.011 () for the all-TDE sample and the subsample with IR flares, respectively. However, the lowest value remains below the 3 confidence level, even though trial factors have not been taken into account yet. In total, at present no significant correlation is found in the population analysis between the ZTF-BTS TDEs and IceCube alert neutrino events. Our work is the first research to quantify the -value of the TDE-neutrino association with actual TDE and neutrino observations.

### v2

@@ 3 Analysis Method
We perform a systematic search in our samples based on the spatial and temporal criteria. The initial step involves examining the spatial correlation between TDEs and alert events. This is realized by cross-matching the coordinates of TDEs with the positions of alert events within their containment radius. In this step, we search for spatially associated events, using the alert events sample with 30 deg2 or the whole alert events sample. According to the of IceCube alert events, the calculation of for alert events is given by . In the next step, we identify the temporal coincident events from the spatially correlated events. The TDEs are required to be discovered before the arrival dates of alert events (i.e., , where is the arrival date of the alert event minus the discovery date of TDE).
To test the correlation between neutrino alerts and TDEs, we define a test statistic (TS). As mentioned in Sec. 2.2, we consider two methods to handle the sample of alert events and calculate the TS value. The first method involves excluding alert events with 30 deg2 (hereafter referred to CUT method). In this method, the number of TDEs that are spatially-temporally associated with the alert events corresponds to the TS values. More specifically, if one TDE is found within the of an alert event and 0, the TS value would be increased by 1. During the calculation of TS values, we allow a single TDE or alert event to contribute repeatedly to the TS values. Specifically, if one alert event is spatially and temporally associated with TDEs, or if one TDE is spatially and temporally associated with alert events, they can each contribute a TS value of (i.e., the total TS will be plus ).
The above CUT method is the simplest approach to determine associations between sources and neutrinos and is widely used (Plavin et al., 2020; Hovatta et al., 2021; Kovalev et al., 2022; Plavin et al., 2023; Suray & Troitsky, 2023; Plavin et al., 2024), but it requires excluding events with large error regions. To make full use of all neutrino alert events, another method does not apply a cut and assigns a weighting factor to each alert (hereafter referred to WEIGHT method). We use a similar formalism as in Kouch et al. (2024) but add an additional Gaussian term to account for the angular distance between the neutrino and the TDE. The weighting factor is expressed as
@@
where , and are the signalness, containment radius estimated from and for the -th alert event, respectively; is the angular distance between TDE and central position of alert event; is the median of of alert events sample. In Eq. (2), the term accounts for the probability that the alert is an astrophysical neutrino rather than a background event. The term considers the location of the TDE within the confidence region of a correlated neutrino; namely, associations that are closer in distance are given higher weights. The term takes into account the fact that an event with a larger error region indicates lower data quality and a higher chance of random association, and thus should be assigned a lower weight. In the WEIGHT method, the TS value is calculated by with given by Eq. (2) when TDEs are found to be spatially-temporally associated with the -th alert event. More specifically, the TS value would be adjusted from 1 in the CUT method to for each TDE associated with the -th alert event in order to reduce the effect of alert events with large in the WEIGHT method.
@@
According to the of the alert events, we estimate that the of our alert events sample is 6.88 deg2. We present the 1-event TS values using the heating map after considering the weight factor (the angular distance term is not included because it depends not only on the event properties) in Fig. 3, where the color bar is the TS value when one TDE is found to be associated with an alert event, the black dashed line indicates the median value of our alert events sample (6.88 deg2), inverted triangles are alert events that are not associated any TDE and the red stars present alert events that are spatially-temporally associated with at least one TDE. One can see that the values of the alert events that are associated with TDEs are larger than the median value.
in the WEIGHT method of the correlation analysis, which depends on the and of the alert event.
Based on these two methods for handling the alert sample and calculating the TS values, we perform Monte Carlo simulations to quantify the correlation between the TDE population and the alert population. Since in our previous work we found that different randomization methods in the MC simulations may have some effect on the results of the correlation analysis (Lu et al., 2024), in the following analysis we will test three different randomization methods.
@@
Method I: Similar to previous works (e.g., Plavin et al., 2020; Hovatta et al., 2021), due to the sensitivity of IceCube only depending on the zenith angle (Aartsen et al., 2017b), we only randomize the RA of the alert events while keeping the DEC, and arrival date fixed to their observed values, and the coordinates of the TDEs also keep unchanged.
@@
Method II: We not only randomize the RA of alert events, as in Method I, but also randomize the positions of the TDEs using the method described in Lu et al. (2024). For each mock TDE source, we randomly select an and a from the observed lists of RA and DEC, then impose an additional random shift to within a region of radius to obtain the position of the mock source. In the simulations, we require the simulated sources not to be outside the sky region of the ZTF-BTS sample, which covers the northern sky with a DEC range from to 90∘ but excluding the Galactic plane region (). We also require the number of mock sources to be the same as the real sample. Finally, we perform KS tests on and distributions to ensure that the mock sources preserve the same distribution pattern of the actual sources.
@@
Method III: The same as Methods I and II, we randomize the RA of alert events. For TDEs, the method used to simulate the mock sample is the same as Buson et al. (2022, 2023); Bellenghi et al. (2023). We generate mock TDEs by randomizing the positions of the sources within a 10∘-radius range from their original positions. During this procedure, we also require that each mock source remains within the sky region of the ZTF-BTS sample. The KS tests are performed in the same manner as in Method II.
@@
By performing 104 MC simulations, we obtain the TS distribution. We can then estimate the chance probability (i.e., the value of rejecting the null hypothesis) according to the fraction of TS values that are equal to or larger than the observed one,
@@
where is the number of TS values equal to or larger than the observed one and is the total number of MC simulations (Davison & Hinkley, 1997).
@@ 4 Results
We first systematically search for associated events that satisfy spatial and temporal criteria. After adopting the CUT method, there are 122 alert events remaining in our sample. Two associated events are found, and all of these events have IR echoes, which are listed in Table 2. Apart from AT2019dsg, the other TDE is associated with one bronze alert events.
Under the WEIGHT method, all 143 alert events are utilized to search for associated events. We found that there are six occurrences of TDEs within the of five alert events. Among them, one TDE (AT2019azh) is located within of two alert events. Additionally, the of two alert events each includes two TDEs. However, we note that the values of these two alert events are significantly large ( 10∘), contributing only negligible TS values. In total, five distinct TDEs are found within of four alert events, two of which have IR echoes, as listed in Table 3.
RAalert | DECalert | Signalness | TDE event | TS1 valueb | IR echoesc | ||||
(deg) | (deg) | (deg) | (days) | ||||||
124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 1 | Yes | 0.022 |
314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1 | Yes | 0.0512 |
@@
c This column indicates whether the TDEs have IR echoes.
RAalert | DECalert | Datealert | Signalness | TDE event | TS1 valuec | IR echoese | ||||
(deg) | (deg) | (deg) | (days) | |||||||
239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020vwl | 1244 | 14.19 | 7.90 | No | 0.0325 |
239.63 | 39.94 | 15.38 | 24/03/07 | 0.606 | AT2020pj | 1526 | 8.72 | 2.68 | No | 0.068 |
127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2022bdw | 522 | 2.31 | 8.78 | No | 0.03782 |
127.18 | 20.74 | 10.16 | 23/07/07 | 0.466 | AT2019azh | 1596 | 4.06 | 6.85 | Yes | 0.022 |
124.54 | 20.74 | 2.39 | 23/02/17 | 0.454 | AT2019azh | 1456 | 2.22 | 2.39 | Yes | 0.022 |
314.08 | 12.94 | 2.95 | 19/10/01 | 0.589 | AT2019dsg | 175 | 1.28 | 9.63 | Yes | 0.0512 |
@@
b The angular distance between alert event and TDE.
c The 1-event TS represents the value that an association event can contribute to the total TS of the correlation.
e This column indicates whether the TDEs have IR echoes.
@@
CUT method | 0.39 | 0.38 | 0.37 | All TDEs |
WEIGHT method | 0.40 | 0.42 | 0.42 | All TDEs |
CUT method | 0.025 | 0.018 | 0.022 | TDEs with IR echoes |
WEIGHT method | 0.038 | 0.044 | 0.055 | TDEs with IR echoes |
By realizing 104 times MC simulations, the results obtained from the three MC methods are summarized in Table 4, which includes the results from both the CUT method and the WEIGHT method. Using the all TDEs sample, the best -values are 0.37 and 0.40 for the CUT method and WEIGHT method, respectively. When using the subsample of TDEs with IR echoes, the best -values are 0.018 and 0.038 for the CUT method and WEIGHT method, corresponding to 2.37 and 2.08, respectively. The most significant correlation with neutrinos is obtained from the subsample of TDEs with IR echoes at a confidence level, which may indicate that TDEs with IR echoes are more likely to correlate with IceCube neutrinos compared to all TDEs sample.
@@ 5.1 The systematic errors of alert events
Although the IceCube collaboration strives to consider all potential errors in estimating the size of the high-energy neutrino error regions, there may still be some unknown systematic errors, such as those arising from ice inhomogeneities, that cannot be completely addressed (e.g. Abbasi et al., 2023b). An upper limit of 1∘ is estimated for this systematic error by Aartsen et al. (2013c). While efforts to constrain this parameter are ongoing (Abbasi et al., 2021b, 2022b), the most recent findings suggest that the associated error is likely negligible (Abbasi et al., 2022b, 2023b).
In this work, we add a systematic error upper limit of 1∘ in quadrature to the four directions of RA and DEC, so that the final uncertainty in each direction is . We then search for associated events among all alert events (having added 1∘ systematic error) and TDEs. However, no additional associated events are found when compared to those listed in Table 3. Therefore, the value is expected to increase as the systematic error grows larger within the range of . When a 1∘ systematic error is added to each alert event, the of the alert events sample is 10.02 deg2. We only adopt the WEIGHT method to test the effect of systematic error on the estimation of values. The values estimated by three MC simulation methods are listed in Table 5, showing an increase of to compared to those in Table 4.
@@
WEIGHT+Syst. | 0.48 | 0.51 | 0.49 | All TDEs |
WEIGHT+Syst. | 0.045 | 0.060 | 0.063 | TDEs with IR echoes |
WEIGHT+Delay | 0.23 | 0.23 | 0.20 | All TDEs |
WEIGHT+Delay | 0.0080 | 0.035 | 0.022 | TDEs with IR echoes |
@@ 5.2 Time delay factor
In this analysis, the temporal coincidence criterion is required that the discovery dates of TDEs are earlier than the arrival dates of alert events. In this scenario, we estimate the values of correlation between TDEs/subsample of ZTF-BTS catalog and IceCube neutrinos, and the most significant correlation is given by the TDEs with IR echoes subsample.
@@
Therefore, we tentatively impose a time delay factor into the weighting scheme to reduce the weight of associated events with large time delays. By combining an additional time delay factor into the weight factor of Eq. 2a, 2b, we reduce the impact of events with large time delays, i.e.
@@
After adding the time delay factor, the TS values are mainly contributed by the AT2019dsg event. By realizing 104 times MC simulations, the values for the three MC simulation methods are summarized in Table 5. Using the WEIGHT method, compared to the results without the time delay factor, the best value of the all-TDEs sample decreases to 0.20, while that for the TDEs with IR echoes subsample decreases to 0.0080 (). After introducing the time delay factor, the subsample with IR echoes also presents the most significant correlation with IceCube neutrinos compared to the all-TDEs sample.
5.3 On the connection between high-energy neutrinos and IR emissions
TDEs have increasingly been discovered through optical transient surveys and have been proposed as potential sources of high-energy neutrinos over the past few years. Interestingly, it is noted that the first three TDE candidates reported to be possibly associated with high-energy neutrinos, including the confirmed TDE AT2019dsg and the candidates AT2019fdr and AT2019aalc (Stein et al., 2021; Reusch et al., 2022; van Velzen et al., 2024), as well as some other similar events reported subsequently (Jiang et al., 2023; Yuan et al., 2024; Li et al., 2024), all exhibited infrared echo emissions. In this work, our analysis also shows that the TDEs with IR flares display a stronger correlation with high-energy neutrinos, which further supports previous findings that the TDEs coinciding with neutrinos tend to have prominent IR echo emissions.
The IR emission typically originates from a dust torus surrounding the supermassive black hole: the the optical-ultraviolet or X-ray radiation from the TDE outburst is absorbed by the dust and re-emitted in the IR band, indicating the presence of a dense environment around such TDEs (Lu et al., 2016; van Velzen et al., 2016; Jiang et al., 2016). Although not fully clear, possible reasons for TDEs with IR echoes are more likely to be associated with IceCube high-energy neutrinos include (Wu et al., 2022; Mou et al., 2021; Winter & Lunardini, 2023; Yuan & Winter, 2023; Yuan et al., 2024) (1) Significant IR echoes imply a dusty/cloudy environment around the black hole, where winds or outflows interacting with the circumnuclear medium can drive the formation of shocks, which accelerate protons to high energies and lead to neutrino production. (2) The dense gas and dust in the dust torus provide target material for the accelerated high-energy protons. Compared to TDEs without a dust torus, such an environment significantly enhances the rate of neutrino production via interactions. (3) The IR photons produced by dust echoes serve as target photons for interactions, and accelerated protons interact with the IR photons, though this requires the protons to be accelerated to sufficiently high energies.
Overall, our analysis, together with the previous works, suggests that the IR echoes are not only an important feature of TDEs, but may also reflect key environmental conditions for neutrino production. Searching for neutrino emission from transients with infrared echoes is therefore a promising direction worth further attention.
@@ 6 Summary
Currently, the main contributors to the IceCube neutrino flux are still obscure. Among various astrophysical sources, TDEs have been suggested as possible emitters of IceCube neutrinos. Therefore, investigating the correlation between the TDE population and IceCube neutrinos could help us better understand whether TDEs could really serve as neutrino emitters. In this work, we study the correlation between the TDEs/TDE candidates from the ZTF-BTS catalog and the population of IceCube neutrinos. Considering that good spatial localization is crucial for establishing reliable association, we employ two methods to handle the alert events with large error circles. One method involves directly excluding alert events with error area 30 deg2, while another method introduces a weighting factor. We quantify the significance of the correlation using MC simulations. We find that the subsample of TDEs with IR echoes presents the most significant correlation with IceCube neutrinos, though still below the confidence level. We also discuss the case that includes a systematic error in the events’ error radii. Given that only the statistical error is publicly available for alert events, we add a systematic error of , which is the upper limit of the systematic error estimated by Aartsen et al. (2013c), in quadrature to the statistical error. There are no additional associated events arising from the inclusion of systematic error and the obtained results could be treated as the upper limits of the values. Furthermore, for the potentially associated events reported so far, the arrival dates of neutrinos are within an interval of year after the discovery dates of the TDEs/TDE candidates. Therefore, we tentatively introduce a time delay () term into the weighting factor in our MC simulation to reduce the weight of large time-delayed events. By combining the time delay factor, the best values decrease to 0.20 and 0.0080 () for the all-TDE sample and the subsample with IR echoes, respectively. However, the lowest value remains below the 3 confidence level, even though trial factors have not been taken into account yet.
In total, at present no significant correlation is found in the population analysis between the ZTF-BTS TDEs and IceCube alert neutrino events. Our work is the first research to quantify the -value of the TDE-neutrino association with actual TDE and neutrino observations. We find that the TDEs with IR flares have a stronger correlation with high-energy neutrinos, consistent with previous findings.
In our sample, only eight TDEs have IR flares and two of them satisfy the spatial-temporal association with high-energy neutrinos. The number of TDEs with IR flares is still limited, which hampers us from obtaining conclusive results. In the future, the persistent optical survey conducted by ZTF-BTS (Perley et al., 2020) is expected to detect more TDEs, in the order of 10 TDEs per year. Moreover, hundreds to thousands of TDEs are expected to be discovered per year by advanced surveys such as the Legacy Survey of Space and Time (LSST) at the Vera Rubin Observatory (Ivezić et al., 2019; Bricman & Gomboc, 2020) and the Wide Field Survey Telescope (WFST, Lin et al. (2022); Wang et al. (2023)). It is also expected that the planned NEO Surveyor (Mainzer et al., 2023), the successor to the NEOWISE mission, will detect more IR flares and help build a larger sample of TDEs with IR echoes, addressing the current issue of too small sample size. By identifying more cases of associations between IR-TDEs and IceCube high-energy neutrinos, we will hopefully further uncover the intrinsic connection between the TDE population and high-energy neutrinos, and understand the role that IR echoes play in the neutrino emission process.